### v1

# Enhancing ECG Classification Robustness with Lightweight Unsupervised Anomaly Detection Filters

Abstract
Continuous electrocardiogram (ECG) monitoring via wearables offers significant potential for early cardiovascular disease (CVD) detection. However, deploying deep learning models for automated analysis in resource-constrained environments faces reliability challenges due to inevitable Out-of-Distribution (OOD) data. OOD inputs, such as unseen pathologies or noise-corrupted signals, often cause erroneous, high-confidence predictions by standard classifiers, compromising patient safety. Existing OOD detection methods either neglect computational constraints or address noise and unseen classes separately. This paper explores Unsupervised Anomaly Detection (UAD) as an independent, upstream filtering mechanism to improve robustness. We benchmark six UAD approaches, including Deep SVDD, reconstruction-based models, Masked Anomaly Detection, normalizing flows, and diffusion models, optimized via Neural Architecture Search (NAS) under strict resource constraints (at most 512k parameters). Evaluation on PTB-XL and BUT QDB datasets assessed detection of OOD CVD classes and signals unsuitable for analysis due to noise. Results show Deep SVDD consistently achieves the best trade-off between detection and efficiency. In a realistic deployment simulation, integrating the optimized Deep SVDD filter with a diagnostic classifier improved accuracy by up to 21 percentage points over a classifier-only baseline. This study demonstrates that optimized UAD filters can safeguard automated ECG analysis, enabling safer, more reliable continuous cardiovascular monitoring on wearables.
I Introduction
Cardiovascular diseases (CVDs) are the leading cause of mortality worldwide, accounting for 19.8 million deaths in 2022 [1]. Early detection and timely intervention, encompassing pharmacological therapy and lifestyle modifications, are critical for improving patient outcomes and mitigating the associated healthcare burden. However, traditional monitoring within clinical settings often fails to capture the sporadic or asymptomatic cardiac events that may serve as early indicators of underlying pathologies. The advent of wearable technology offers a promising avenue for continuous, ambulatory electrocardiogram (ECG) monitoring, facilitating the detection of transient abnormalities without impeding patients’ daily activities. The large-scale deployment of continuous ECG monitoring generates vast quantities of data that preclude manual analysis, necessitating automated interpretation. Over the past decade, machine learning (ML), and particularly deep learning (DL), has demonstrated remarkable success in analyzing complex ECG signals [2].
Despite these advancements, the transition from controlled datasets to real-world clinical deployment, especially on resource-constrained wearable devices, poses significant challenges. Beyond the fundamental requirements of high accuracy and computational efficiency, a critical challenge arises from the inevitability of encountering Out-of-Distribution (OOD) data. In the context of ambulatory ECG analysis, OOD data typically manifests in two critical forms: ECG signals exhibiting pathologies unseen during the model’s training (unseen CVD classes), and signals that are so severely corrupted by noise artifacts (e.g., electrode motion or muscle activity) that they are rendered unsuitable for analysis. Standard classifiers are prone to making erroneous, often high-confidence, predictions on such inputs, which can undermine the reliability of the system [3, 4].
Addressing this challenge through purely supervised learning is impractical, as it requires an exhaustive, labeled dataset encompassing all possible CVDs and noise profiles encountered in daily life. Consequently, research has focused on methods for OOD detection and Uncertainty Quantification (UQ). Existing approaches often rely on analyzing the outputs or latent features of supervised classifiers (e.g., using MC Dropout, Deep Ensembles, post-hoc energy scores, or specialized architectures) [5, 6, 7, 8]. While these methods are valuable, they frequently address the detection of unseen classes and the identification of noise in isolation. Furthermore, many sophisticated approaches overlook the stringent computational constraints required for real-time deployment on wearable devices.
To address these gaps, this paper investigates the application of Unsupervised Anomaly Detection (UAD) as an independent filtering mechanism deployed upstream of a diagnostic classifier, as illustrated in Figure 1. UAD methods learn a compact representation of the in-distribution data without requiring labeled examples of anomalies, enabling the detection of significant deviations caused by either novel pathologies or noise. We conduct a systematic evaluation of diverse UAD methodologies, including Deep Support Vector Data Description (Deep SVDD), reconstruction-based methods (autoencoders (AE) and variational autoencoders (VAE)), Masked Anomaly Detection (MAD), Denoising Diffusion Probabilistic Models (DDPM), and normalizing flows (NF). Our analysis focuses specifically on the trade-off between detection performance and computational efficiency essential for deployment in resource-constrained environments.
Figure 1: Schematic overview of the proposed robust ECG monitoring architecture. An independent UAD filter is deployed upstream of the diagnostic classifier. The filter assesses incoming ECG signals. If they are deemed anomalous (OOD or unsuitable for analysis due to noise), the signal is rejected, thereby protecting the classifier from unreliable inputs and improving overall system robustness.
The main contributions of this work are summarized as follows:
  • We benchmark six distinct UAD methodologies, optimized via Neural Architecture Search (NAS) under strict resource constraints ($\leq$512k parameters). We evaluate their efficacy for the detection of OOD CVD classes and signals unsuitable for analysis due to noise using the PTB-XL and BUT QDB datasets.
  • We empirically identify Deep SVDD as the superior methodology for this application, demonstrating that it offers the optimal balance between computational efficiency and detection performance
  • We validate the efficacy of an integrated classifier-filter system in a realistic deployment simulation incorporating unseen CVD classes and calibrated real-world ambulatory noise. We show that the upstream UAD filter significantly enhances diagnostic robustness, yielding an accuracy improvement of up to 21.0 percentage points over a classifier-only baseline.
The remainder of this paper is organized as follows: Section II reviews related work. Section III details the datasets, the investigated UAD methodologies, and the experimental setup. Section IV presents the empirical results. Section V discusses the findings. Finally, Section VI concludes the paper.
II Related Work
The deployment of ML models for ECG analysis in real-world clinical settings introduces significant challenges beyond controlled environments. Key among these are generalization across different data distributions (domain shift), robustness to varying signal quality, and the critical safety requirement of identifying inputs that fall outside the model’s training, known as OOD detection or open-set recognition.
II-A Domain Generalization and Distribution Shifts
A significant challenge in ML-based ECG analysis is performance degradation due to distribution shifts (e.g., variations in demographics or hardware). This is typically addressed as a domain generalization problem, aiming to maintain performance on known classes despite these shifts. Ballas and Diou [9] leveraged intermediate features from a ResNet-18 to improve generalization for 12-lead ECG classification across hospital databases. Similarly, Soltanieh et al. [10] demonstrated that Self-Supervised Learning (SSL) techniques (SimCLR, BYOL, SwAV) can learn representations that generalize well for arrhythmia detection across different OOD datasets. While these works address robustness to domain shifts, they do not specifically tackle the detection of entirely unseen disease classes or the identification of data unsuitable for analysis due to noise.
II-B Signal Quality Detection and Uncertainty Quantification
The reliability of automated ECG analysis is heavily dependent on input signal quality, as noise artifacts (e.g., muscle activity, electrode motion) prevalent in ambulatory monitoring can severely impair diagnostic accuracy. The literature addresses this through Signal Quality Assessment (SQA) and Uncertainty Quantification (UQ). Traditional SQA approaches focus on the explicit detection and classification of noise contamination. Satija et al. [11] proposed a framework using modified ensemble empirical mode decomposition (CEEMD) and temporal features to categorize ECG signals by noise type, facilitating targeted denoising or false alarm reduction. Alternatively, UQ methods address signal quality indirectly by identifying when a diagnostic model’s confidence is compromised due to impaired input. Jahmunah et al. [12] developed a Dirichlet DenseNet model to quantify uncertainty in myocardial infarction diagnosis. Utilizing predictive entropy, they demonstrated the model’s ability to convey low confidence when analyzing noisy signals, aiming to mitigate the risk of erroneous predictions based on corrupted data.
In contrast, our work employs UAD as an independent, upstream mechanism to explicitly filter out signals deemed entirely unsuitable for analysis due to noise, irrespective of the specific noise type, thereby preventing unreliable data from reaching the classifier.
II-C Unsupervised Anomaly Detection in ECG
UAD methods identify deviations from a learned distribution of normal data without labeled anomalies. In ECG analysis, UAD has been explored primarily to differentiate abnormal patterns from normal sinus rhythm or to enhance supervised learning. Atamny et al. [13] conducted a comparative study of various unsupervised models, including AEs, VAEs, diffusion models, NFs, and Gaussian Mixture Models (GMMs). Trained solely on normal data, these models utilized reconstruction errors or likelihood estimations to differentiate abnormal ECG signals from healthy ones, with VAEs demonstrating the highest performance. In a different approach aimed at addressing the challenges of highly imbalanced (long-tail) datasets, Jiang et al. [14] employed self-supervised anomaly detection as a pretraining mechanism. They utilized masking and restoration techniques with multi-scale cross-attention to enhance feature representation, thereby improving the subsequent classification of rare cardiac anomalies.
While these works utilize UAD methodologies, their objectives differ significantly from ours. They focus respectively on binary diagnostic separation (healthy vs. abnormal) or improving the recognition of rare, but known, classes. In contrast, our work concentrates on deploying UAD as an independent, upstream filter designed to reject entirely unseen disease classes and signals rendered unsuitable for further analysis by noise.
II-D Open-Set Recognition in ECG Applications
Open-set recognition, the detection of unknown inputs during deployment, is also prevalent in biometric authentication using cardiac signals. Studies in this domain focus on rejecting unregistered subjects. Dong et al. [15] proposed a robust authentication system using multi-modal pretraining and specialized loss functions to enforce clear decision boundaries. Wu et al. [16] utilized compressed CNNs and user-specific feature vectors for efficient open-set identification. Although these studies address the open-set problem, the application domain involves fundamentally different objectives and data characteristics compared to clinical diagnosis.
II-E Detection of Unseen Cardiovascular Diseases
The literature most closely related to our research involves detecting OOD samples corresponding to unseen CVDs, primarily through UQ techniques or post-hoc analysis of supervised classifiers. Barandas et al. [5] evaluated UQ methods (Deep Ensembles, MC Dropout) in a multi-label ECG setting. They assessed OOD detection using the PTB-XL dataset by treating specific superclasses (myocardial infarction and hypertrophy) as OOD, similar to our design. However, their focus was on UQ applied to supervised classifiers rather than independent UAD filters, and they did not explicitly address the simultaneous detection of signals unsuitable for analysis due to noise. Elul et al. [6] employed a multi-head architecture where an unknown class is inferred if all binary classifiers output a negative prediction, supplemented by MC Dropout for uncertainty estimation. While addressing OOD detection, their strategy is tied to the classifier’s architecture, not a dedicated UAD filter, and overlooks efficiency constraints crucial for wearables. Yu et al. [7] proposed a trustworthy diagnosis method using post-hoc Energy and ReAct techniques on a CNN-Attention classifier to recognize OOD heart diseases. Eidheim [8] utilized latent space features from a trained supervised classifier (e.g., xResNet1D101) and applied traditional anomaly detection methods (e.g., Mahalanobis Distance) to these features. Similar to UQ approaches, these methods rely on analyzing the supervised model’s behavior or features.
In summary, while prior work has made significant strides in OOD detection and UQ for ECG analysis, these efforts often rely on supervised features, address unseen classes and noise in isolation, or overlook the efficiency constraints required for resource-limited environments. Our work addresses these gaps by systematically evaluating and optimizing a diverse range of UAD methods for the simultaneous detection of both unseen CVDs and ECG samples unsuitable for analysis due to noise, focusing specifically on the performance-efficiency trade-off essential for deployment on wearable devices.
III Methodology
This section details the datasets utilized, the UAD methods investigated, and the experimental protocols designed to evaluate their performance in detecting both OOD CVDs and ECG signals unsuitable for analysis due to noise.
III-A Datasets and Preprocessing
We used three datasets in this study. PTB-XL is a large, publicly accessible clinical ECG database comprising 21,799 12-lead ECG recordings, each 10 seconds in duration, from 18,869 patients in its most recent version 1.0.3 [17, 18, 19]. The recordings are annotated according to the SCP-ECG standard, covering diagnostic, form, and rhythm statements provided by cardiologists. Diagnostic labels are organized hierarchically into five superclasses: Normal (NORM), Myocardial Infarction (MI), Conduction Disturbance (CD), ST/T Change (STTC), and Hypertrophy (HYP). There are potentially multiple superclasses per record, as they are not mutually exclusive. The dataset facilitates standardized evaluation through pre-defined, patient-aware, 10-fold cross-validation splits, stratified by diagnosis, sex, and age. ECG signals are provided at both 500 Hz and 100 Hz sampling frequencies. We chose the 500 Hz samples.
The Brno University of Technology ECG Quality Database (BUT QDB) is designed for evaluating signal quality algorithms [20, 19]. It contains 18 long-term, single-lead ECG recordings (sampled at 1,000 Hz) collected during normal daily activities. Expert annotations classify signal quality into three categories: Class 1 (clear visibility of QRS complex, P waves and T waves), Class 2 (only QRS complexes reliably detectable), and Class 3 (signal unsuitable for analysis, QRS complexes undetectable).
The MIT-BIH Noise Stress Test database provides realistic noise recordings characteristic of ambulatory ECGs [21, 19]. It includes three half-hour recordings representing Baseline Wander (bw), Muscle Artifact (ma), and Electrode Motion Artifact (em). The ’em’ record, utilized in this study for noise injection, contains significant electrode motion artifacts along with baseline wander and muscle noise.
All ECG recordings were segmented into 10-second windows. These windows were resampled to a uniform length of 512 timesteps, a configuration determined during pre-screening tests to optimize the balance between computational efficiency and model performance. For the PTB-XL dataset, z-score normalization was applied, as it yielded superior performance. For the BUT QDB dataset, instance normalization proved more effective.
III-B Unsupervised Anomaly Detection Methods
We investigated six UAD methods for detecting anomalous ECG signals. To optimize hyperparameters and evaluate the performance-efficiency trade-off, a NAS using random search (100 trials) was conducted for each method. Models were constrained to a maximum of 512k parameters to ensure suitability for resource-constrained devices.
Deep SVDD aims to map normal data into a compact hypersphere of minimum volume [22]. The training objective minimizes the distance of representations from a predefined center, enforcing constraints (no bias terms, unbounded activations, fixed center) to prevent hypersphere collapse. The anomaly score is the squared Euclidean distance to the center in the output space. A 1D ResNet architecture was employed. The NAS optimized the number of layers, filters, kernel sizes, strides, and the latent dimension size.
MAD utilizes a self-supervised approach similar to masked autoencoders [23]. During training, 5% of time steps (determined via pre-screening) are masked with random values, and the model reconstructs the original values using bidirectional context. The anomaly score is the sum of reconstruction errors calculated by sequentially masking and reconstructing each time step during inference. A standard transformer architecture with an initial patching layer was used. The NAS varied the number of transformer blocks, hidden dimensions, attention heads, linear layer sizes, patch size, and patch overlap.
Reconstruction-based methods (AE and VAE) assume that models trained solely on normal data will fail to accurately reconstruct anomalous inputs [24]. We utilized the L2 reconstruction error as the anomaly score. The VAE differs by employing a probabilistic encoder regularized by a KL divergence loss toward a standard normal distribution [25]. Stacked 1D convolution layers were used for both encoder and decoder architectures. The NAS optimized the number of layers, filters, kernel sizes, strides, and latent dimension size. For the VAE, the KL loss weight was also varied.
DDPMs involve a fixed forward diffusion process that gradually adds Gaussian noise over T timesteps, and a learned reverse denoising process [26]. For anomaly detection, we adopted the approach of Wyatt et al. [27]: an anomalous input is partially diffused and then reconstructed. The model, trained only on the normal data manifold, attempts to ”repair” abnormal regions during denoising. A 1D U-Net architecture with convolution and attention layers was implemented using the denoising-diffusion-pytorch package [28]. The NAS varied the diffusion steps, U-Net architecture parameters (layers, filters, attention heads, hidden dimensions), and the U-Net objective (predicting noise, the original input or a variable from angular parametrization as in Salimans and Ho [29]).
NFs are exact-likelihood generative models that learn complex data distributions by transforming a simple base distribution through invertible mappings [30, 31]. We implemented a multiscale GLOW architecture [32] with affine coupling layers and invertible 1x1 convolutions, utilizing the normflows package [33]. The exact probability density of a sample serves as the anomaly score. Stacked 1D convolutions were used for the subnetwork within the coupling layers. The NAS optimized the subnetwork architecture (layers, filters, kernel sizes) and the flow architecture (number of coupling blocks, split/squeeze dimensions and ratios).
Figure 2: Visualization of the realistic noise injection process applied to the PTB-XL dataset. A clean ECG signal (bottom left) is corrupted by adding calibrated noise (top center), sourced from the MIT-BIH noise stress test database. The resulting noisy ECG signal (bottom right) simulates real-world ambulatory conditions where the signal quality is unsuitable for analysis.
III-C Experimental Setup
Two main experiments were conducted to assess the UAD methods and the overall system performance. During the NAS phase (Experiment 1), each configuration was trained three times to account for initialization variability. For final testing (Experiments 1 and 2), evaluations were repeated five times, reporting the mean and standard deviation of the relevant metrics.
Experiment 1 - Detection Performance of OOD CVD Classes and Signals Unsuitable for Analysis due to Noise
This experiment evaluated the performance of the UAD methods across two distinct tasks: Detection of ODD classes and of signals unsuitable for analysis due to noise. Using the PTB-XL dataset, four scenarios were created. In each scenario, one diagnostic superclass (MI, CD, STTC, or HYP) was designated as the OOD class. All samples containing the OOD class, including those where it co-occurred with in-distribution classes, were removed from the training data (folds 1-8). The validation (fold 9) and test (fold 10) sets included both in-distribution and OOD samples. Performance was measured using the Area Under the Receiver Operating Characteristic curve (AUC). Using the BUT QDB dataset, the data was split patient-wise to avoid data leakage. The training set included patients with only Class 1 and Class 2 noise labels (acceptable quality). The test set included patients exhibiting Class 3 noise (unsuitable for analysis). A 10-second window was labeled as unsuitable for analysis due to noise (anomalous) if at least 1 second within that window was annotated as Class 3.
Figure 3: Performance-efficiency trade-offs (Pareto fronts) resulting from the NAS in Experiment 1 across five anomaly detection scenarios. Each panel illustrates the trade-off between detection performance, measured by AUC (higher is better), and computational efficiency, measured by parameter count (lower is better). The upper-left region represents the optimal trade-off. Deep SVDD, NF and MAD consistently demonstrate superior efficiency compared to AE, VAE and DDPM in OOD CVD detection tasks. All methods achieve very high AUC for the detection of ECG signals unsuitable for analysis due to noise (rightmost panel).
Experiment 2 - Integrated Classifier-Filter System Performance
While Experiment 1 benchmarked the intrinsic detection performance of the UAD methods, it does not quantify the impact of such a filter on the overall diagnostic pipeline. The second experiment was therefore designed to assess the practical benefit of employing a UAD filter upstream of a standard multilabel classifier in a realistic deployment scenario involving both OOD classes and signals unsuitable for analysis due to noise.
A high-performing multilabel classification model was required for this evaluation. We selected the resnet1d_wang model, identified as the best-performing architecture on the PTB-XL diagnostic superclasses in the comprehensive benchmark by Strodthoff et al. [34]. This model achieves a macro AUC of 0.930, comparable to more complex state-of-the-art methods like L5G-Net (0.9357) and X-ECGNet (0.936). Their PyTorch implementation [35] was ported to TensorFlow. The classifier was trained in the same four scenarios as Experiment 1, using only the four in-distribution classes and excluding the OOD class from training, validation, and initial testing sets to assess in-distribution performance. Deep SVDD was selected as the filter mechanism, as it demonstrated the best average performance across both detection of OOD classes and of signals unsuitable for analysis due to noise in Experiment 1. The hyperparameter configuration yielding the top result for each specific scenario was utilized.
To simulate realistic conditions, the PTB-XL validation and test sets were modified to include calibrated amounts of noise using the nst script from PhysioNet [36]. The ’em’ noise record from the MIT-BIH Noise Stress Test Database was injected (see Figure 2). The Signal-to-Noise Ratio (SNR) used by nst is defined based on the peak-to-peak amplitude of QRS complexes. To ensure validity of the simulation, it was crucial to determine the appropriate SNR for injection by estimating the noise level present in real-world signals unsuitable for analysis (BUT QDB Class 3 samples). The estimation required a specialized wavelet denoising procedure designed to preserve R-peak morphology for the subsequent amplitude calculation. We employed the ’db6’ mother wavelet due to its morphological similarity to the QRS complex, good energy compaction and good time localization. Using a decomposition level of 11 (appropriate for the 1,000 Hz sampling rate), Donoho’s universal threshold [37, 38] was applied with soft thresholding. The noise standard deviation was estimated from the median absolute deviation of the first-level detail coefficient to ensure accurate noise variance estimation. This calibrated noise was applied to approximately 16.5% of the validation and test data, matching the prevalence of Class 3 noise windows (signals unsuitable for analysis) observed in our BUT QDB test set. Crucially, this injection ratio was applied across both the in-distribution and OOD samples within each scenario (HYP, CD, STTC, MI). This stratification ensures a sufficient representation of samples that are both unsuitable for analysis due to noise and belong to an unseen class, allowing for a robust evaluation of the system’s ability to handle these concurrent challenges.
A custom accuracy metric was defined to evaluate the combined system, accounting for both filtering and classification efficacy. An outcome was counted as correct only if:
  1. 1.
    An ECG sample that should be rejected (OOD or too noisy) was successfully rejected by the filter.
  2. 2.
    An ECG sample that should be accepted was accepted and correctly classified by the classifier.
All other outcomes were considered incorrect. The performance of the classifier-only approach and the integrated classifier-filter system were compared using this custom accuracy. The impact of the filter was further analyzed by plotting the custom accuracy against the rejection rate.
IV Results
This section presents the empirical results of the experiments designed to evaluate the efficacy of various UAD methods for identifying OOD CVDs and ECG signals unsuitable for analysis due to noise, and the subsequent impact of integrating the best-performing method into a classification pipeline.
IV-A Experiment 1 - Detection Performance of OOD CVD Classes and Signals Unsuitable for Analysis due to Noise
The first experiment involved a comprehensive NAS to optimize six UAD methods (Deep SVDD, MAD, AE, VAE, NF, and DDPM) across five distinct anomaly detection scenarios. Four scenarios involved identifying an unseen OOD diagnostic superclass (HYP, CD, STTC, or MI), and one scenario focused on detecting ECG signals unsuitable for analysis due to noise. The objective was to maximize detection performance (AUC) while maintaining computational efficiency, constrained by a 512k parameter limit.
Figure 3 illustrates the Pareto fronts resulting from the NAS, depicting the trade-off between AUC and parameter count for each method and scenario. A general trend across the OOD scenarios (HYP, CD, STTC, MI) is that performance initially increases with model complexity before reaching a plateau, often well within the parameter constraint. The results reveal significant differences in performance profiles among the UAD methods. Deep SVDD, NF, and MAD consistently demonstrated superior efficiency, generally occupying the upper-left regions of the Pareto fronts. This indicates higher AUC scores with fewer parameters compared to traditional reconstruction-based approaches (AE, VAE) and DDPM, which were often dominated across the range of parameter counts in OOD tasks.
Figure 4: System performance (custom accuracy) as a function of the rejection rate across the four OOD scenarios. The rejection rate (x-axis) is controlled by varying the decision threshold of the Deep SVDD filter. A stricter threshold yields a higher rejection rate. Dashed horizontal lines represent the baseline accuracy of the classifier-only system (0% rejection rate). The curves demonstrate a critical trade-off. Accuracy initially increases as anomalies are correctly filtered out, peaks at an optimal operating point and subsequently declines as the filter becomes overly aggressive and rejects in-distribution data.
The difficulty of the detection task also varied significantly depending on the scenario. Detecting HYP as OOD was the most tractable OOD task, with two methods achieving AUC scores above 0.80. Conversely, detecting STTC and MI proved more challenging, with peak AUC scores generally remaining below 0.70. In contrast, the detection of ECG signals unsuitable for analysis due to noise (NOISE panel in Figure 3) was highly successful across all methods. Nearly all evaluated architectures achieved AUC scores approaching 1.0, irrespective of the parameter count, indicating that severely corrupted ECG signals are easily distinguishable from acceptable quality signals by these UAD approaches.
TABLE I summarizes the best-performing AUC results achieved during the NAS for each UAD method and scenario. Deep SVDD demonstrated the most robust and consistent high performance across the varied OOD tasks. It achieved the highest AUC in three of the four OOD scenarios: CD (0.769 $\pm$ 0.006), STTC (0.678 $\pm$ 0.015), and MI (0.692 $\pm$ 0.012). In the HYP scenario, MAD (0.811 $\pm$ 0.001) and NF (0.811 $\pm$ 0.0003) achieved the top result, marginally outperforming Deep SVDD (0.798 $\pm$ 0.004). Traditional reconstruction-based methods (AE, VAE) and DDPM generally underperformed in OOD detection; for example, in the MI scenario, the AUC for DDPM was only 0.518 $\pm$ 0.009.
For the detection of signals unsuitable for analysis due to noise, all methods performed exceptionally well, with AUCs exceeding 0.95. Deep SVDD (0.992 $\pm$ 0.002) and AE (0.991 $\pm$ 0.001) yielded the best results. Given its superior average performance across both detection of OOD classes and of signals unsuitable for analysis due to noise, Deep SVDD was selected as the filter mechanism for the subsequent experiment.
IV-B Experiment 2 - Integrated Classifier-Filter System Performance
The second experiment evaluated the practical utility of deploying a UAD filter (using the optimized Deep SVDD model) upstream of a standard multilabel classifier in a simulated deployment environment containing both OOD classes and injected realistic noise. Performance was evaluated using a custom accuracy metric, which balances the need to filter anomalies with the need to process valid data. This metric credits the system only if an anomalous sample is rejected or if an in-distribution sample is accepted and correctly classified.
Figure 4 illustrates the relationship between the system’s rejection rate and the resulting custom accuracy. The rejection rate (x-axis) is varied by adjusting the anomaly score threshold of the Deep SVDD filter. A stricter threshold increases the filter’s sensitivity, thereby increasing the rejection rate. The dashed horizontal lines indicate the baseline accuracy of the classifier-only system, corresponding to a 0% rejection rate. In all scenarios, as the rejection rate increases from zero, the custom accuracy rises sharply, significantly surpassing the baseline performance. This improvement is attributed to the successful filtering of OOD and noisy samples that the classifier would otherwise misclassify. However, the curves demonstrate a clear trade-off between robustness and utility. The accuracy peaks at an optimal rejection rate, typically between 20% and 40% across the different scenarios. In this optimal region, the benefit of correctly rejecting anomalies significantly outweighs the cost of occasionally rejecting in-distribution data. Beyond this optimum, the accuracy declines because the filter becomes overly aggressive, increasingly rejecting in-distribution samples that the classifier could have correctly processed, thus diminishing the overall utility of the system.
TABLE II compares the custom accuracy of the baseline classifier-only system with the integrated classifier-filter system operating at the optimal rejection rate. The baseline classifier exhibited relatively low custom accuracy, ranging from 0.460 (MI) to 0.509 (HYP). This highlights the significant negative impact of OOD and noisy samples on overall system reliability in a realistic setting. The introduction of the Deep SVDD filter substantially improved the custom accuracy across all scenarios. The largest improvement was observed in the CD scenario, where accuracy increased from 0.466 $\pm$ 0.006 to 0.676 $\pm$ 0.006, representing an improvement of 21.0 percentage points. Similar significant percentage point gains were observed for MI (+18.1), STTC (+15.6), and HYP (+15.4). These results demonstrate the efficacy of the proposed filtering approach in enhancing the robustness of ECG classification systems.
V Discussion
TABLE I: Summary of the best detection performance (AUC) achieved by each UAD method during NAS (Experiment 1), constrained to 512k parameters. Results are reported as mean $\pm$ standard deviation over five runs. For each scenario, the top-performing method and any method within one standard deviation are highlighted in bold font.
UAD
Method
AUC
HYP CD STTC MI Noise
Deep SVDD 0.798 $\pm$ 0.004 0.769 $\pm$ 0.006 0.678 $\pm$ 0.015 0.692 $\pm$ 0.012 0.992 $\pm$ 0.002
MAD 0.811 $\pm$ 0.001 0.749 $\pm$ 0.002 0.645 $\pm$ 0.001 0.665 $\pm$ 0.003 0.971 $\pm$ 0.002
[…]
VAE 0.781 $\pm$ 0.001 0.652 $\pm$ 0.001 0.617 $\pm$ 0.002 0.579 $\pm$ 0.001 0.984 $\pm$ 0.002
NF 0.811 $\pm$ 0.0003 0.723 $\pm$ 0.0003 0.666 $\pm$ 0.003 0.650 $\pm$ 0.004 0.975 $\pm$ 0.001
DDPM 0.764 $\pm$ 0.001 0.607 $\pm$ 0.003 0.576 $\pm$ 0.001 0.518 $\pm$ 0.009 0.957 $\pm$ 0.001
TABLE II: Comparison of overall system robustness (custom accuracy) between the baseline (classifier only) and the integrated system (classifier + optimized Deep SVDD filter) in the realistic deployment simulation (Experiment 2). Results are reported as mean $\pm$ standard deviation. The integrated system performance is reported at the optimal rejection rate, i.e. the peak accuracy observed in Figure4. The superior approach in each scenario is highlighted in bold font.
UAD
Method
Custom Accuracy
HYP CD STTC MI
Classifier only
0.509
$\pm$ 0.007
0.466
$\pm$ 0.006
0.462
$\pm$ 0.002
0.460
$\pm$ 0.006
Classifier + Filter
0.663
$\pm$ 0.011
0.676
$\pm$ 0.006
0.618
$\pm$ 0.004
0.641
$\pm$ 0.007
This study systematically investigated the application of UAD as an independent filtering mechanism to enhance the robustness of ECG classification systems in resource-constrained environments. The objective was to identify inputs unsuitable for reliable analysis, specifically unseen CVD classes and signals unsuitable for analysis due to noise. Through a comprehensive benchmark utilizing NAS under strict computational constraints ($\leq$ 512k parameters), we demonstrated that an upstream UAD filter significantly improves the overall accuracy and reliability of a downstream classifier, achieving an improvement in custom accuracy of up to 21.0 percentage points.
V-A Interpretation of UAD Method Performance
The comprehensive benchmark (Experiment 1) revealed significant disparities in the efficacy of the investigated UAD methodologies. Deep SVDD demonstrated the most robust overall performance, achieving the highest AUC in the CD (0.769), STTC (0.678), and MI (0.692) scenarios. Its superiority likely stems from its geometric, one-class objective, mapping in-distribution data into a minimal-volume hypersphere, which bypasses weaknesses inherent in other methods.
Traditional reconstruction methods (AE and VAE) assume anomalous data is harder to reconstruct than normal data; however, recent work suggests this assumption is often violated [39], potentially explaining their lower performance in detecting subtle OOD CVDs. AE and VAE performed relatively well for Hypertrophy (HYP) (AUCs $\approx$ 0.78). HYP is often characterized by significantly increased QRS amplitudes [40], which likely fall far outside the training distribution—a scenario AEs are generally effective at detecting.
MAD demonstrated strong performance, particularly for HYP (AUC 0.811). By reframing reconstruction as a contextual infilling task, MAD is sensitive to contextual anomalies involving violations of local temporal dependencies. This aligns well with detecting secondary repolarization abnormalities (”strain patterns”) sometimes seen in HYP [40]. Although theoretically suited for STTC (also a contextual anomaly), MAD’s performance was moderate (AUC 0.645). This may be due to its potential weakness in detecting widespread anomalies; if the bidirectional context is also anomalous, MAD might successfully predict a masked point based on its correlated, anomalous neighbors, resulting in a missed detection.
NFs, performing explicit density estimation, also excelled in the HYP scenario (AUC 0.811), likely detecting the distributional shift caused by altered voltage characteristics. However, their performance for MI and CD was modest. NFs are susceptible to assigning high likelihoods to OOD data [41]. Furthermore, in high-dimensional data, likelihood calculation can be dominated by local correlations [42, 43], causing the model to deem a sequence probable based on plausible local wave shapes while failing to recognize anomalous global structures.
DDPMs exhibited the weakest performance across OOD CVD tasks. We utilized standard Gaussian noise, following Ho et al. [26]. However, since ECG power is concentrated in lower frequencies (0-50Hz) [44], Gaussian noise (with uniform spectral density) may insufficiently corrupt these low-frequency components [27]. Consequently, low-frequency ECG anomalies might survive the partial noising process and be preserved during denoising, resulting in a low reconstruction error and poor detection performance, even though high frequencies (>>70Hz) also carry clinical value [44].
In contrast to the varied performance in OOD CVD detection, all methods demonstrated excellent efficacy (AUC $\geq$ 0.957) in detecting ECG signals unsuitable for analysis due to noise. Severe noise artifacts manifest as gross, high-amplitude distortions that represent global point anomalies far outside the distribution of acceptable quality ECGs, readily detected by all evaluated UAD paradigms.
V-B Variability in OOD Detection Difficulty
The results in TABLE I reveal significant variability in detecting different diagnostic superclasses as OOD. HYP was the most tractable (best AUC $\approx$ 0.81), followed by CD (best AUC $\approx$ 0.77). Detecting MI and STTC proved substantially more challenging (best AUCs $\approx$ 0.69 and 0.68). This disparity is attributable to the distinctiveness of their ECG manifestations and the effects of clinical confounding.
The relative ease of detecting HYP and CD likely stems from their pronounced impact on the QRS complex (ventricular depolarization). HYP typically manifests as augmented QRS amplitudes [45, 46, 47]. CD, particularly bundle branch blocks (BBBs), involves significant prolongation of the QRS duration and fundamental alterations in QRS morphology [48]. These substantial changes in amplitude, duration, and wave shape create features qualitatively distinct from other superclasses, facilitating clearer separation in the feature space.
In contrast, STTC and MI primarily involve abnormalities in the ventricular repolarization phase (ST segment and T-wave) [49]. These changes are often more subtle and variable than depolarization changes. Furthermore, there is inherent clinical overlap, as MI manifests primarily through STTC features (e.g., ST elevation/depression) [49].
However, the primary driver of the difficulty in isolating STTC and MI as OOD is likely clinical confounding, where in-distribution pathologies mimic the OOD class. Repolarization abnormalities are not exclusive to STTC or MI; they frequently occur as secondary effects of HYP and CD. HYP can induce ”strain” patterns (downsloping ST depression and T-wave inversion) [40] that overlap with changes seen in STTC and MI. Likewise, conditions in CD, such as LBBB, can produce significant secondary ST-T changes [48]. In the context of UAD, this confounding is critical. When STTC or MI is held out, the training data still contains samples with HYP and CD. Consequently, the UAD model learns that the morphological features associated with HYP strain and secondary CD changes are part of the in-distribution manifold. When subsequently presented with true ischemic patterns (STTC/MI), the model struggles to identify them as anomalous because similar morphologies are already represented within its learned boundary of normality, directly accounting for the lower AUC scores observed.
V-C Clinical Implications and System Robustness
The integration of the Deep SVDD filter upstream of the 1D ResNet classifier demonstrated substantial improvements in system robustness (Experiment 2). In the presence of both OOD classes and calibrated real-world noise, the baseline classifier exhibited poor custom accuracy (0.460 to 0.509). This highlights the vulnerability of standard diagnostic models when deployed outside controlled environments, where they are prone to making erroneous, often high-confidence, predictions on inputs they were not trained to handle. The UAD filter significantly mitigated this vulnerability, yielding percentage points improvements in custom accuracy ranging from 15.4 (HYP) to 21.0 (CD).
This enhancement is critical for the safe deployment of automated ECG analysis, particularly on wearable devices where continuous monitoring increases the probability of encountering novel pathologies or severe noise artifacts. By rejecting inputs that cannot be reliably processed, the system inherently adheres to the precautionary principle fundamental to clinical practice. This mechanism significantly reduces the risk of compromised patient safety by mitigating the likelihood that erroneous, high-confidence predictions on OOD data could precipitate inappropriate or potentially harmful clinical decisions.
The analysis of the rejection rate versus custom accuracy (Figure 4) reveals a crucial trade-off. While increasing the rejection rate initially improves accuracy by filtering out anomalies, an overly aggressive threshold leads to the rejection of in-distribution samples, diminishing the system’s utility. The optimal operating point, observed between 20% and 40% rejection in this study, must be carefully calibrated based on the specific clinical application.
The magnitude of the observed improvement is naturally dependent on the prevalence of OOD and noisy samples in the test set. However, our experimental design aimed to provide a realistic estimation. We designated only one diagnostic superclass as OOD and injected noise into 16.5% of the samples, mirroring the prevalence of signals unsuitable for analysis due to noise (Class 3) observed in the BUT QDB ambulatory dataset. Furthermore, many commercially available wearable devices like smartwatches are trained to detect a narrower set of conditions (e.g., Atrial Fibrillation [50]). In such applications, the spectrum of potential OOD CVDs is considerably larger, suggesting that the positive impact of an upstream UAD filter might be even more pronounced than demonstrated in this study.
V-D Comparison with Related Work
This work distinguishes itself from prior research that primarily focuses on Uncertainty Quantification (UQ) [5, 11] or post-hoc analysis of supervised classifiers [7, 8] for OOD detection. While those methods are valuable, they inherently rely on the features learned by the supervised model. In contrast, our approach employs independent UAD models trained solely on in-distribution data, offering a modular filtering mechanism applicable to any downstream classifier. Furthermore, this study addresses the simultaneous detection of both unseen pathologies and ECG signals unsuitable for analysis due to noise under strict efficiency constraints, a combination often overlooked in the literature. The utilization of NAS to systematically optimize UAD architectures specifically for the performance-efficiency trade-off is crucial for deployment in resource-constrained ECG analysis.
V-E Limitations and Future Work
This study has several limitations. First, the evaluation of computational efficiency relied on parameter count as a proxy for resource usage. Practical deployment on wearable hardware requires direct measurement of power consumption, latency, and memory footprint. Future work will involve implementing the Pareto-optimal models on representative microcontroller units (MCUs) to quantify these metrics and explore optimization techniques such as quantization and pruning.
Second, the methodology for simulating realistic noise in Experiment 2 relies on estimating the SNR of BUT QDB Class 3 samples. Determining the true SNR of severely corrupted ambulatory signals is inherently challenging, introducing the possibility of miscalibration in the simulation. However, the impact on the study’s conclusions is likely minimal, given the very high efficacy (AUC $\geq$ 0.957) of the UAD methods in detecting these severe noise artifacts (TABLE I).
Third, the resampling of ECG windows to 512 timesteps (51.2 Hz) was an empirically driven trade-off necessary to manage the substantial computational demands of the extensive NAS process. Although diagnostic information contained in higher frequencies may have been attenuated, our pre-screening tests confirmed that this resolution led to no loss in performance compared to the original frequencies.
Fourth, the study utilized ECG as the sole modality. Incorporating contextual sensor modalities, such as accelerometer data, could improve the ability to distinguish between signals unsuitable for analysis due to noise and those representing unseen CVD classes. This differentiation would enable differential system responses (e.g., discarding a sample unsuitable for anaylsis due to noise vs. alerting a clinician about an unseen CVD).
Finally, opportunities exist to further enhance UAD performance. Future research could explore advanced techniques, such as training NFs on high-level semantic representations rather than raw input data to mitigate issues with local correlations [43], utilizing simplex noise instead of Gaussian noise for DDPMs to better corrupt low-frequency components [27], or combining complementary UAD methods within an ensemble framework.
VI Conclusion
In this paper, we addressed the critical challenge of ensuring the reliability of machine learning models for ECG analysis when deployed in real-world settings, where encountering unseen cardiovascular diseases and signals unsuitable for analysis due to noise is inevitable. We conducted a comprehensive benchmark and comparative analysis of six UAD methods, optimized via NAS under strict resource constraints suitable for wearable devices.
Our evaluation demonstrated that Deep SVDD offers the superior balance of efficiency and detection performance for identifying both OOD pathologies and severe noise artifacts, outperforming traditional reconstruction-based methods, MAD, NFs, and DDPMs in this application.
By integrating the optimized Deep SVDD model as an upstream filter to a standard diagnostic classifier, we achieved a significant enhancement in overall system robustness. In a realistic deployment simulation incorporating both OOD classes and calibrated ambulatory noise, the integrated system yielded improvements in accuracy of up to 21.0 percentage points compared to the baseline classifier-only approach.
This work underscores the critical role of independent UAD mechanisms in safeguarding the performance of automated ECG analysis systems, paving the way for safer and more clinically reliable, low-resource wearable ECG monitoring systems at scale. Future research will focus on hardware implementation, direct power consumption analysis, and the exploration of advanced techniques to incorporate context information and further improve detection capabilities.
Code Availability
The code supporting the findings of this study will be made publicly available upon publication.